OpenStudy (anonymous):

x+1/5 = 2x-2/6 please solve

OpenStudy (arbershabani97):

x-2x=-2/6-1/5 -x=-10/30-6/30 -x=-16/30 X=?

OpenStudy (anonymous):

I assume its (x+1)/5=(2x-2)/6 then multiply 30 on both sides 6(x+1)=5(2x-2) distribute: 6x+6=10x-10 add 10 both sides: 6x+16=10x subtract 6x both: 16=4x divide 4 both sides: x=4

OpenStudy (arbershabani97):

probably, but we have to wait till we get the exact question from zoe000999

OpenStudy (anonymous):

This is the question:Which of the following equations have the same solution? Solve all three equations separately and explain each step. I. 3x + 2 = 5x – 6 II. 2x – 2 = 8 III. the quantity of x plus 1, divided by 5 equals the quantity of 2x minus 2, divided by 6 I already got 1 and 2 but I just needed help with 3 :) @arbershabani97 @jamesumner93

OpenStudy (arbershabani97):

then it is (x+1)/5=(2x-2)/6 and 10x-10=6x+6 4x=16 x=4

OpenStudy (anonymous):

Thank you soooo much omg @arbershabani97 xx

OpenStudy (arbershabani97):

you're welcome